Doubly‑charged calcium ions, Ca2+, are released into your bloodstream and flow through your capillaries. The diameter of a capillary is about 9.00 μm, calcium ion concentrations are approximately 1.00 mol/m3, and typical electric current flow due to the Ca2+ ions is around 1.2 nA. Calculate the drift speed ?drift of the calcium ions.
Answer #1

Number of moles

n=1*6.02*1023 =6.02*1023 moles

charge

Q=2e =2*1.6*10-19=3.2*10-19 C

Area

A=pi*D2/4 =pi*(9*10-6)2/4 =6.36*10-11 m2

Drift speed

Vd =I/nAQ =(1.2*10-9)/[(6.02*1023 )(6.36*10-11)(3.2*10-19 )]

Vd=9.8*10-5 m/s